### v1

@@ C.2 Experimental Setting
Our analysis is based on T2I-CompBench++ (Huang et al., 2025), which provides curated prompts across attributes (color, shape, texture), spatial relations (2D and 3D), non-spatial relations, complex prompts, and numeracy. Each prompt is paired with images from multiple text-to-image models and annotated with human evaluation scores. All resources (prompts, images, and scores) come from the benchmark; our contribution is to analyze how evaluation metrics align with these annotations using outputs from SD v1.4, SD v2, Structured Diffusion (Feng et al., 2023b), Composable Diffusion (Liu et al., 2022), Attend-and-Excite (Chefer et al., 2023a), and GORS (Huang et al., 2023).
We evaluate five embedding-based metrics (PickScore (Kirstain et al., 2023), CLIPScore (Hessel et al., 2021), HPS (Wu et al., 2023), ImageReward (Xu et al., 2024), BLIP-2 (Li et al., 2023a)), two image-only metrics (CLIP-IQA (Wang et al., 2023), Aesthetic Score (Schuhmann et al., 2022)), and five VQA-based metrics (B-VQA (Huang et al., 2023), DA Score (Singh & Zheng, 2023), TIFA (Hu et al., 2023), DSG (Cho et al., 2023), VQA Score (Lin et al., 2024b)), covering embedding similarity, perceptual quality, and VQA-style reasoning.
@@ C.3 Correlation Analysis of Evaluation Metrics
Several broader insights emerge from these results. First, no single metric achieves strong and consistent correlation across all compositional categories, indicating that reliance on a single signal is insufficient. Second, despite its widespread use (Rombach et al., 2022; Nichol et al., 2021; Ruiz et al., 2023; Brooks et al., 2023; Kumari et al., 2023; Kang et al., 2023a; Chefer et al., 2023b; Podell et al., 2023; Chen et al., 2023; Li et al., 2023b; Nguyen & Tran, 2024), CLIP never ranks among the top metrics, underscoring its limitations as a standalone measure. Third, embedding-based metrics, particularly ImageReward and HPS, frequently appear among the strongest. Fourth, while VQA-based metrics are competitive, they are not uniformly superior and are occasionally outperformed by embedding-based approaches. Finally, image-only metrics such as CLIP-IQA and Aesthetic remain consistently weak, as expected since they do not assess text–image alignment.
@@ Appendix E Effect of Multi-Clip on Multi-Backward Optimization
Appendix F Human Evaluation Protocol
@@
Appendix G Quality and Diversity Evaluation
@@
Appendix H Time and Memory Usage Analysis
We also evaluate the computational efficiency of CARINOX by measuring runtime and VRAM usage on three backbones: PixArt-$\alpha$, SD-Turbo, and SDXL-Turbo. Both CARINOX and ReNO introduce additional overhead compared to the raw backbones, but differ in their requirements due to the complexity of their optimization pipelines.
As shown in Table 9, CARINOX requires more resources than ReNO, reflecting its integration of multiple reward models and iterative optimization. For instance, on SD-Turbo, VRAM usage increases from $15$,GB with ReNO to $33$,GB with CARINOX, while runtime rises from $20$,s to $60$,s. Similar patterns are observed on PixArt-$\alpha$ and SDXL-Turbo.
@@ Appendix H Time and Memory Usage Analysis
Appendix I Pseudo code for Noise Optimization and Exploration
@@
1:$p$ (prompt), $G_{\theta}$ (One-Step T2I Model), $S_{1\ldots N}$ (random seeds), $\mathcal{R}_{1\ldots M}$ (reward functions), $T$ (iterations), $\eta$ (learning rate), $\tau$ (grad clip), $\lambda_{\text{reg}}$ (regularization strength)
@@
Appendix J Alternative Benchmark: GenEval
To further validate our method, we evaluate on the GenEval benchmark, which measures compositional alignment across categories such as single/multi-object generation, counting, color attribution, and spatial positioning. Results in Table 10 show that CARINOX achieves competitive or superior performance across both SD-Turbo and SDXL-Turbo backbones. In particular, it delivers strong improvements in color attribution and overall mean scores, matching or surpassing state-of-the-art baselines including ReNO and large-scale commercial systems. These findings confirm that CARINOX generalizes well beyond the primary benchmarks used in the main paper.
@@
Appendix K Additional Qualitative Examples

### v2

@@ C.2 Experimental Setting
Our analysis is based on T2I-CompBench++ (Huang et al., 2025), which provides curated prompts across attributes (color, shape, texture), spatial relations (2D and 3D), non-spatial relations, complex prompts, and numeracy. Each prompt is paired with images from multiple text-to-image models and annotated with human evaluation scores. All resources (prompts, images, and scores) come from the benchmark; our contribution is to analyze how evaluation metrics align with these annotations using outputs from SD v1.4, SD v2, Structured Diffusion (Feng et al., 2023b), Composable Diffusion (Liu et al., 2022), Attend-and-Excite (Chefer et al., 2023), and GORS (Huang et al., 2023).
We evaluate five embedding-based metrics (PickScore (Kirstain et al., 2023), CLIPScore (Hessel et al., 2021), HPS (Wu et al., 2023), ImageReward (Xu et al., 2024), BLIP-2 (Li et al., 2023a)), two image-only metrics (CLIP-IQA (Wang et al., 2023), Aesthetic Score (Schuhmann et al., 2022)), and five VQA-based metrics (B-VQA (Huang et al., 2023), DA Score (Singh and Zheng, 2023), TIFA (Hu et al., 2023), DSG (Cho et al., 2023), VQA Score (Lin et al., 2024b)), covering embedding similarity, perceptual quality, and VQA-style reasoning.
@@ C.3 Correlation Analysis of Evaluation Metrics
Several broader insights emerge from these results. First, no single metric achieves strong and consistent correlation across all compositional categories, indicating that reliance on a single signal is insufficient. Second, despite its widespread use (Rombach et al., 2022; Nichol et al., 2021; Ruiz et al., 2023; Brooks et al., 2023; Kumari et al., 2023; Kang et al., 2023a; Chefer et al., 2023; Podell et al., 2023; Chen et al., 2023; Li et al., 2023b; Nguyen and Tran, 2024), CLIP never ranks among the top metrics, underscoring its limitations as a standalone measure. Third, embedding-based metrics, particularly ImageReward and HPS, frequently appear among the strongest. Fourth, while VQA-based metrics are competitive, they are not uniformly superior and are occasionally outperformed by embedding-based approaches. Finally, image-only metrics such as CLIP-IQA and Aesthetic remain consistently weak, as expected since they do not assess text–image alignment.
@@ Appendix E Effect of Multi-Clip on Multi-Backward Optimization
To provide a quantitative view of why Multi-Clip is needed, Table 8 reports both the mean and standard deviation of the per-reward gradient norms across iterations. Two patterns are consistent: (i) the scales remain mismatched throughout the trajectory (e.g., VQA stays the largest even at the end: Iter 50 mean 4.03 vs. HPS 0.22 and DA 0.16), so aggregation can remain dominated by a subset of rewards beyond the first few steps; and (ii) the standard deviations are large—especially early on (e.g., Iter 1: VQA std 112.12, ImageReward std 72.67)—indicating high variability and occasional extreme gradients that can abruptly steer the update direction. Multi-Clip addresses both by clipping each reward gradient independently before aggregation, preventing persistent or sporadic domination by any single reward.
Iter VQA Score ImageReward HPS DA Score
Mean Std Mean Std Mean Std Mean Std
1 39.09 112.12 11.52 72.67 2.37 3.36 6.27 15.19
10 15.25 71.11 2.79 11.84 0.77 1.08 1.37 3.28
20 8.70 31.43 1.36 4.71 0.45 0.57 0.64 1.70
30 6.36 30.54 0.82 5.85 0.32 0.39 0.33 0.92
40 4.85 14.26 0.51 1.77 0.26 0.29 0.21 0.54
50 4.03 10.64 0.39 1.89 0.22 0.29 0.16 0.53
Table 8: Mean and standard deviation of per-reward gradient norms $\lVert\nabla_{\bm{\epsilon}}\mathcal{R}\rVert_{2}$ at selected optimization iterations (averaged over prompts/seeds used in the Multi-Clip analysis).
Appendix F Adaptive Reward Weighting (Ablation)
We include the following experiment as an ablation to assess whether replacing uniform reward weights ($\lambda_{i}{=}1$) with category-aware correlation-based weights improves optimization.
While a fixed combination of reward models provides a stable optimization framework, we evaluate a variant that assigns weights to each reward function based on its measured correlation with human preferences within each T2I-CompBench++ compositional category. The weight for each reward $\mathcal{R}_{i}$ is computed as:
$w_{i}=\frac{4\rho_{i}}{\sum_{j}\rho_{j}},$ (15)
where $\rho_{i}$ denotes the correlation coefficient of $\mathcal{R}_{i}$ with human evaluations for the corresponding category (Table˜5). The weights are normalized to sum to $4$ to keep the overall reward scale comparable to the uniform-weight setting. In this ablation, we use the ground-truth category labels provided by T2I-CompBench++ to select the category-specific weights.
(1) + CARINO $0.8519$ $0.7336$ $0.8043$ $0.2437$ $0.3920$ $0.5903$ $0.3269$ $0.4906$ $0.5542$
(1) + CARINO (AW) $0.8608$ $0.7430$ $0.8060$ $0.2411$ $0.4083$ $0.6036$ $0.3239$ $0.4896$ $0.5595$
(2) + CARINO (AW) $0.8437$ $0.7238$ $0.8042$ $0.2767$ $0.3969$ $0.5848$ $0.3138$ $0.4877$ $0.5539$
(3) + CARINO (AW) $0.8215$ $0.7562$ $0.8015$ $0.2770$ $0.4030$ $0.6098$ $0.3150$ $0.4813$ $0.5582$
Table 9: Adaptive reward weighting (AW) ablation on T2I-CompBench++. AW replaces uniform weights with category-specific correlation-based weights, using the benchmark’s ground-truth category labels to select weights for each prompt.
Table 9 shows that adaptive weighting yields only marginal changes relative to uniform weighting. On SD-Turbo, AW slightly improves the mean score (0.5542 $\rightarrow$ 0.5595), while on SDXL-Turbo it is essentially unchanged (0.5554 $\rightarrow$ 0.5539), and on PixArt-$\alpha$ DMD it is again marginal (0.5560 $\rightarrow$ 0.5582). Overall, these results suggest that the uniform weighting used in the main method is already a strong and robust default, and that correlation-based category-aware reweighting provides limited additional benefit under our current setup.
Appendix G Human Evaluation Protocol
@@
Appendix H Quality and Diversity Evaluation
@@
Appendix I Time and Memory Usage Analysis
I.1 Measured Runtime and VRAM
We evaluate the computational efficiency of CARINOX by measuring runtime and VRAM usage on three backbones: PixArt-$\alpha$, SD-Turbo, and SDXL-Turbo. All measurements are obtained on an NVIDIA H100 GPU.
@@ I.1 Measured Runtime and VRAM
As shown in Table 11, both CARINOX and ReNO introduce additional overhead compared to the raw backbones, but differ in their requirements due to the complexity of their optimization pipelines. CARINOX incurs higher runtime and VRAM usage than ReNO due to its multi-reward optimization and best-of-$N$ exploration. For example, on SD-Turbo, VRAM increases from $15$ GB (ReNO) to $33$ GB (CARINOX), and runtime increases from $20$ s to $60$ s. Similar trends hold on PixArt-$\alpha$ and SDXL-Turbo.
The additional overhead in CARINOX stems from two factors: (i) multi-reward optimization, where each iteration evaluates multiple reward models and backpropagates their gradients, and (ii) best-of-$N$ exploration, where optimization is repeated independently for multiple noise seeds and the best result is selected. In particular, some of our rewards rely on heavier vision-language or VQA-style models, which are more demanding in both VRAM and runtime; using them jointly amplifies the per-iteration cost. Importantly, CARINOX is modular with respect to both components: if future work provides reward models with lower memory/latency or comparable quality at reduced cost, our framework can directly benefit by swapping in these scorers without changing the optimization procedure. Similarly, advances in seed exploration (e.g., more sample-efficient initialization search that achieves the same best-of-$N$ gains with fewer candidates) would reduce the number of optimized seeds required, lowering total runtime proportionally.
I.2 NFE-Matched Comparison
To address concerns about unequal inference-time compute, we additionally report NFE-matched comparisons, where NFE denotes the number of generator evaluations. For multi-step diffusion models, we set $\mathrm{NFE}=K$ where $K$ is the number of denoising steps (we use $K=50$ throughout). For single-step backbones (SD-Turbo, SDXL-Turbo, and PixArt-$\alpha$ DMD), a single forward generation corresponds to $\mathrm{NFE}=1$. Iterative noise-optimization methods (ReNO and CARINO) therefore use $\mathrm{NFE}=T$, where $T$ is the number of optimization iterations. For CARINOX, the total compute scales as $\mathrm{NFE}=N\times T$, where $N$ is the number of explored seeds and $T$ is the number of optimization iterations per seed. Under a fixed budget of $\mathrm{NFE}=50$, we report CARINO (50 iterations) and a compute-matched CARINOX configuration with $N=2$ and $T=25$ ($2\times 25$). We also include our main CARINOX setting ($N=5$, $T=50$, $\mathrm{NFE}=250$) to illustrate the gains from additional inference-time scaling beyond the fixed-budget regime.
Model NFE Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ 2D Spatial $\uparrow$ 3D Spatial $\uparrow$ Numeracy $\uparrow$ Non-Spatial $\uparrow$ Complex $\uparrow$ Mean $\uparrow$
SD v1.4 50 0.3765 0.3576 0.4156 0.1246 0.3030 0.4461 0.3079 0.3080 0.3299
SD v2.1 50 0.5065 0.4221 0.4922 0.1342 0.3230 0.4579 0.3127 0.3386 0.3734
SDXL 50 0.5879 0.4687 0.5299 0.2133 0.3566 0.4988 0.3119 0.3237 0.4114
PixArt-$\alpha$-ft 50 0.6690 0.4927 0.6477 0.2064 0.3901 0.5058 0.3197 0.3433 0.4468
Structured + SD v2.1 50 0.4990 0.4218 0.4900 0.1386 0.3224 0.4550 0.3111 0.3355 0.3717
Attn-Exct + SD v2.1 50 0.6400 0.4517 0.5963 0.1455 0.3222 0.4550 0.3111 0.3355 0.4072
InitNO + SD v2.1 50 0.7038 0.4694 0.5212 0.2027 0.3524 0.4892 0.3105 0.3574 0.4258
(1) SD-Turbo 1 0.5252 0.4434 0.4888 0.1881 0.3112 0.4914 0.3095 0.3349 0.3866
(1) + ReNO 50 0.7800 0.6200 0.7500 0.2200 0.3800 0.5700 0.3200 0.4800 0.5150
(1) + CARINO (T=50T{=}50) 50 0.8519 0.7336 0.8043 0.2437 0.3920 0.5903 0.3269 0.4906 0.5542
(1) + CARINOX (N=2N{=}2, T=25T{=}25) 50 0.8383 0.7216 0.7940 0.2413 0.4015 0.6011 0.3264 0.4867 0.5531
(1) + CARINOX (N=5N{=}5, T=50T{=}50) 250 0.8633 0.7609 0.8229 0.2588 0.4155 0.6248 0.3372 0.5041 0.5734
(2) SDXL-Turbo 1 0.5959 0.4038 0.5472 0.2303 0.3612 0.4863 0.3114 0.3430 0.4099
(2) + ReNO 50 0.7800 0.6000 0.7400 0.2600 0.3900 0.5600 0.3100 0.4700 0.5137
(2) + CARINO (T=50T{=}50) 50 0.8492 0.7203 0.7977 0.2858 0.4069 0.5835 0.3141 0.4859 0.5554
(2) + CARINOX (N=2N{=}2, T=25T{=}25) 50 0.8430 0.7141 0.7838 0.2951 0.4076 0.5803 0.3138 0.4880 0.5531
(2) + CARINOX (N=5N{=}5, T=50T{=}50) 250 0.8697 0.7482 0.8270 0.3010 0.4117 0.5992 0.3232 0.4922 0.5715
(3) PixArt-$\alpha$ DMD 1 0.4145 0.3487 0.3667 0.2213 0.3441 0.4896 0.3061 0.3466 0.3547
(3) + ReNO 50 0.6400 0.5700 0.7200 0.2500 0.3900 0.5600 0.3100 0.4600 0.4875
(3) + CARINO (T=50T{=}50) 50 0.8260 0.7528 0.7967 0.2620 0.4031 0.6144 0.3146 0.4782 0.5560
(3) + CARINOX (N=2N{=}2, T=25T{=}25) 50 0.8175 0.7463 0.7803 0.2989 0.4019 0.5960 0.3139 0.4697 0.5522
(3) + CARINOX (N=5N{=}5, T=50T{=}50) 250 0.8545 0.7721 0.8076 0.3272 0.4164 0.6295 0.3256 0.4878 0.5776
Table 12: NFE-matched evaluation on T2I-CompBench++. For multi-step diffusion models and their inference-time variants, we report results at $K=50$ denoising steps ($\mathrm{NFE}=50$). For single-step backbones, we report the raw generation reference ($\mathrm{NFE}=1$), a fixed compute budget setting ($\mathrm{NFE}=50$), and our main inference-time scaling setting ($\mathrm{NFE}=250$).
Table 12 indicates that the benefits of our approach are not tied to using a larger compute budget. In the $\mathrm{NFE}=50$ regime, CARINO and CARINOX (N=2N{=}2, T=25T{=}25) are the top-performing entries among the methods reported: they outperform the multi-step diffusion backbones (SD v1.4/v2.1/SDXL/PixArt-$\alpha$-ft), the inference-time attention methods on SD v2.1 (Structured Diffusion and Attend-and-Excite), the multi-step noise optimizer InitNO, and the single-step baseline ReNO. This shows that even when we match the evaluation budget to the standard 50-step setting, optimizing the single-step generator with our reward combination yields stronger compositional alignment than both multi-step generation and prior inference-time optimization strategies.
Appendix J Effect of Norm-Guided Regularization (Ablation)
We include this experiment as an ablation to isolate the impact of the norm-guided regularization term used in CARINOX. In this setting, we remove the regularization term from the optimization objective and keep all other components unchanged (same rewards, Multi-Clip, learning rate, number of iterations, and seed selection protocol).
Figure 8 visualizes the optimization trajectory at several iterations. Without regularization (top rows), the optimized samples progressively exhibit high-frequency noise, color saturation, and noticeable artifacts, indicating a distributional shift away from the generator’s latent prior. This degradation becomes more pronounced at later iterations, even when the prompt-level attributes or entities appear partially satisfied. In contrast, the full CARINOX setting (bottom rows) preserves visual realism throughout optimization while still improving compositional alignment, suggesting that the norm constraint plays a practical role in stabilizing optimization and preventing drift into unrealistic regions of the image manifold.
Figure 8: Ablation of the norm-guided regularization term. We show intermediate results across optimization iterations for two prompts. Without regularization (top rows), optimization increasingly drifts toward noisy and artifact-prone outputs (distributional shift). With regularization (bottom rows), optimization remains visually stable while improving alignment.
Appendix K Pseudo code for Noise Optimization and Exploration
@@ I.2 NFE-Matched Comparison
1:$p$ (prompt), $G_{\theta}$ (One-Step T2I Model), $S_{1\ldots N}$ (random seeds), $\mathcal{R}_{1\ldots M}$ (reward functions), $K(\cdot)$ (noise regularizer), $T$ (iterations), $\eta$ (learning rate), $\tau$ (grad clip), $\gamma$ (regularization strength)
@@
Appendix L Alternative Benchmark: GenEval
To further validate our method, we evaluate on the GenEval benchmark, which measures compositional alignment across categories such as single/multi-object generation, counting, color attribution, and spatial positioning. Results in Table 13 show that CARINOX achieves competitive or superior performance across both SD-Turbo and SDXL-Turbo backbones. In particular, it delivers strong improvements in color attribution and overall mean scores, matching or surpassing SOTA baselines including ReNO and large-scale commercial systems. These findings confirm that CARINOX generalizes well beyond the primary benchmarks used in the main paper.
@@
Appendix M Additional Qualitative Examples